### v1

# The Design Space of Lockfiles Across Package Managers

Abstract.
Software developers reuse third-party packages that are hosted in package registries. At build time, a package manager resolves and fetches the direct and indirect dependencies of a project. Most package managers also generate a lockfile, which records the exact set of resolved dependency versions. Lockfiles are used to reduce build times; to verify the integrity of resolved packages; and to support build reproducibility across environments and time. Despite these beneficial features, developers often struggle with their maintenance, usage, and interpretation. In this study, we unveil the major challenges related to lockfiles, such that future researchers and engineers can address them.
We perform the first comprehensive study of lockfiles across 7 popular package managers, npm, pnpm, Cargo, Poetry, Pipenv, Gradle, and Go. First, we highlight how the content and functions of lockfiles differ across package managers and ecosystems. Next, we conduct a qualitative analysis based on semi-structured interviews with 15 developers. We capture first-hand insights about the benefits that developers perceive in lockfiles, as well as the challenges they face to manage these files. Following these observations, we make 4 recommendations to further improve lockfiles, for a better developer experience.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0
1. Introduction
Software reuse is an undeniable best practice in software engineering (Krueger, 1992). Software developers reuse third-party libraries for commodity features such as logging, cryptography, arithmetic, or I/O. Reuse helps developers save time, pool quality assurance efforts, and focus on the core, application specific features of their project (Mohagheghi and Conradi, 2007). Yet, software reuse at scale introduces major engineering challenges (Cox, 2019). The use of third-party libraries, including the transitive ones, often results in large dependency trees. Managing these increasingly complex dependency trees is hard because developers have to resolve compatible versions across the whole tree (Rausch et al., 2017), stay on top of updates to patch bugs and vulnerabilities (Ladisa et al., 2023), and ensure that new changes do not break the build (Frank et al., 2024).
Package managers are tools that assist developers in the management of reused libraries (Spinellis, 2012). They automate and simplify the resolution, download, configuration, installation, and update of dependencies. When using a package manager, developers define the dependencies of a project in a dependency specification file. Then, the package manager determines compatible versions of these dependencies and all indirect dependencies, fetches them from a package registry, and installs them to further build the project. During this process, some package managers save the exact list of resolved dependencies and their versions into a special file: the lockfile (Yu et al., 2024).
Lockfiles are widely used throughout the software development lifecycle, supporting reproducible builds (Lamb and Zacchiroli, 2022a), defending the software supply chain (Liu et al., 2022), and contributing to improved build performance (Aïdasso et al., 2025). However, lockfiles also introduce additional complexity, and their maintenance requirements are not fully understood, a subject hardly touched by software engineering research. Lockfiles have been mentioned as security best practices (Kabir et al., 2022), important for software supply chain security (Williams et al., 2025), related to Software Bill of Materials (SBOM) generation (Bi et al., 2024), and notable for new dependency management tools (Cleare and Iacob, 2018). However, none of these works provide an in-depth analysis of lockfiles, why they exist and what they should contain. To our knowledge, there is no systematic analysis and comparison of the different strategies in the field to generate and evolve lockfiles. There is no grounded evidence on how developers handle and make the best out of their lockfiles.
In this study, we address this critical gap in the software engineering literature by conducting a comprehensive analysis of lockfiles across 7 package managers. These package managers serve the top programming languages of 2024 (Cass, 2024): npm and pnpm for Javascript, Cargo for Rust, Poetry and Pipenv for Python, Gradle for Java, and Go for Golang. We analyze the official documentation and implementation of these package managers, and compare the strategies they implement to generate and maintain lockfiles. We then quantitatively evaluate lockfile usage in the wild, by collecting GitHub projects and their lockfile commit rates. To complement this analysis, we conduct interviews with 15 developers with diverse backgrounds and experiences. They contribute to active open source projects spanning various domains such as networking, gaming, CLI, and APIs, and use the package managers under study. Through these interviews, we gather valuable first-hand insights into the developers’ motivations for using or avoiding lockfiles, as well as the challenges they face when working with them.
Our paper is the first study that highlights how the functionalities and usage of lockfiles differ across package managers and ecosystems, resulting in varying levels of developer satisfaction. While some approaches, such as the Go strategy for lockfiles, encourage strong developer engagement, others, such as Gradle’s approach, fail to address key software engineering challenges. Our findings clearly show that a simple, principled approach to the lockfile lifecycle leads to a better developer experience. We also gather evidence of the challenges that developers face with lockfiles, such as limited human readability, difficulty locking indirect dependencies, and delays in receiving dependency updates. Overall, developers generally trust lockfiles and view them as a valuable tool for improving the efficiency and security of dependency management. We conclude our study with 4 recommendations for developers and package manager maintainers to encourage the adoption of lockfiles in all projects.
All data and interview protocols used in this study are publicly available: https://github.com/chains-project/lockfiles-comprehensive-study. In summary, our key contributions are,
  • a comprehensive comparison of lockfiles across 7 top package managers, with a deep analysis of the key dimensions of the lockfile problem space.
  • a qualitative assessment of the benefits and challenges of lockfiles, through 15 developer interviews, demonstrating that developers well acquainted with lockfiles see essential value for addressing important software engineering challenges.
  • a regimen of 4 recommendations to further improve lockfiles for dependency management, to inform future research in the field as well as engineering of the future generation of package managers.
2. Definitions and Context
Software developers who wish to reuse third-party packages, typically specify the configurations for these packages in a dependency specification file. Then, developers can rely on a package manager to help them install and maintain reusable code packages efficiently. When developers invoke an action that involves resolving the dependencies, the package manager parses the dependency specification file, queries the default or specified alternative registries, determines the required package versions, and downloads them. During this process, it may also record the resolved direct and indirect package versions in a lockfile. Developers can review this file to inspect the exact versions selected by the package manager, and may commit it to version control for reproducible builds, improved security, and efficiency. In subsequent project builds the package manager can refer to the lockfile to quickly resolve required dependency versions, perform integrity checks against the downloaded packages and ensure deterministic project builds. This process is illustrated in Figure 1. In the remainder of this section, we define these different key concepts of the package manager, which will be used throughout the paper.
2.1. Definitions
Figure 1. A canonical workflow of interactions between a developer, the dependency specification file, the package manager, the lockfile, and the package registry.
Definition 0.
A package manager facilitates code reuse by ensuring that all the third-party packages required by the project, directly or indirectly, are downloaded with the correct versions. Most popular programming languages have one or more package managers (e.g. Maven for Java or Cargo for Rust). Package managers typically provide developers with a command line interface and sometimes a graphical user interface to interact with the package registry.
Definition 0.
A package registry is as a central platform for sharing code across multiple projects. Depending on the registry, developers can upload packaged binaries (e.g., JAR files of Java bytecode) or packaged source code, allowing others to download the code to be reused within their own projects. Popular programming languages have official package registries: npm.com is widely used for JavaScript, Crates.io for Rust, PyPI.org for Python, etc.
Definition 0.
A dependency specification is a file that specifies the packages that a software project directly depends on, i.e., it specifies the direct dependencies of the project. For each dependency, the developers can either declare a specific version, a practice commonly known as pinning, or specify a range of versions, or specify no version which usually means allow resolving the ‘latest compatible‘ version. The package manager takes the dependency specification file as input to build the project. The dependency specification file also includes additional information related to dependency management, such as metadata about the project or exclusion of certain dependencies. The dependency specification file is primarily maintained by developers, and also by automated dependency bots such as Dependabot.
Definition 0.
A lockfile is a file that is automatically generated by the package manager and that is not meant to be edited by developers. The lockfile records the resolved versions of all direct and indirect dependencies that need to be included in a project. Some package managers may include extra details, such as the hash of the dependency specification file content or the package manager version. Package managers use lockfiles when (re)building projects, such as for fetching the exact versions of dependencies required, and verifying their integrity. Lockfiles may also be read by developers during code reviews to audit resolved dependencies, and updated by automated dependency bots during dependency updates.
Table 1. The 7 package managers analyzed in this study, including their associated programming languages, registries, manifests, and lockfiles.
Package manager Language Package registry Dependency specification file Lockfile
npm CLI JavaScript npm.com package.json package-lock.json
pnpm JavaScript npm.com package.json pnpm-lock.yaml
Cargo Rust crates.io Cargo.toml Cargo.lock
Poetry Python PyPi.org pyproject.toml poetry.lock
Pipenv Python PyPi.org Pipfile Pipfile.lock
Gradle Java central.sonatype.com build.gradle gradle.lockfile
Go Go github.com go.mod go.mod & go.sum
2.2. Considered Package Managers
We select 7 popular package managers for our study based on the popularity of programming languages in 2024 (Cass, 2024). From the ten most popular languages, we identify those with widely used package managers that support lockfiles: JavaScript, TypeScript, Python, Java, and Go. We add Rust due to its significant upward trend. Table 1 outlines the 7 package managers analyzed in this study, along with their associated language, registries, dependency specification files, and lockfiles.
For instance, npm CLI is the default package manager for JavaScript, and the npm registry is the default package registry. Library developers can upload their JavaScript code to the npmjs.com registry. The dependency specification file and lockfile for npm CLI are in JSON format and are named package.json and package-lock.json, respectively. Excerpts from a package.json and its corresponding package-lock.json are given in 1 and 2, respectively.
In the case of Go, there is no centralized package registry. Instead, Go modules are downloaded from version control registries such as GitHub. Module versions are tracked using an index that feeds available new versions. The integrity verification of Go packages is done based on a checksum database. Moreover, in Go, the go.mod file works both as the dependency specification file and the lockfile. An additional file, named go.sum, is generated by the package manager, to record the checksums of the downloaded modules. The go.mod file is not meant to be edited manually by developers but should be modified using specific commands. 3 and 4 show an example go.mod file and a go.sum file.
Table 2 provides the list of selected package manager versions, including the number of lines of code in the specified version, and the number of available releases up to that version. The release counts are obtained from the official releases page of each package manager, except for pnpm, where the release count is based on the information from the npm registry.
Table 2. Details about the exact versions of the 7 package managers under study.
Package manager Analyzed Version Lines of code Released versions
npm CLI npm/cli (11.1.0) 459k (JavaScript) 557
pnpm pnpm/pnpm (10.1.0) 112k (TypeScript) 1103
Cargo rust-lang/cargo (0.85.0 ) 210k (Rust) 119
Poetry python-poetry/poetry (2.0.1) 46k (Python) 88
Pipenv pypa/pipenv (2024.4.1) 140k (Python) 396
Gradle gradle/gradle (8.12.1) 784k (Groovy) 531k (Java) 367k (Kotlin) 148
Go golang/go (1.24.0) 2052k (Go) 243
1{
2 "name": "@asyncapi/modelina",
3 "version": "3.8.0",
4 "description": "Library for generating data models based on inputs such as AsyncAPI, OpenAPI, or JSON Schema documents",
5...
6"dependencies": {
7 "@apidevtools/json-schema-ref-parser": "^11.1.0",
8 "@apidevtools/swagger-parser": "^10.1.0",
9 "@asyncapi/multi-parser": "^2.1.1",
10 "@asyncapi/parser": "^3.1.0",
11...
Listing 1: An excerpt from the dependency specification file (package.json) of the project asyncapi/modelina
1{
2 "name": "@asyncapi/modelina",
3 "version": "3.8.0",
4 "lockfileVersion": 3,
5...
6"node_modules/@apidevtools/json-schema-ref-parser": {
7 "version": "11.1.1",
8 "resolved": "https://registry.npmjs.org/@apidevtools/...",
9 "integrity": "sha512-WZyqzPLrWctQB8qB8uGivvzWgvtcVqlmR...",
10 "dependencies": {
11 "@jsdevtools/ono": "^7.1.3",
12...
Listing 2: An excerpt from the lockfile (package-lock.json) of the project asyncapi/modelina
1module github.com/piplabs/story
2
3go 1.22.11
4
5require (
6 cosmossdk.io/api v0.7.5
7 // NOTE(https://github.com/cosmos/rosetta/issues/76): Rosetta requires cosmossdk.io/core v0.12.0 erroneously ...
8 cosmossdk.io/core v0.12.0
9 cosmossdk.io/depinject v1.0.0
10 cosmossdk.io/errors v1.0.1 // indirect
11...
Listing 3: An excerpt from the dependency specification file (go.mod) of the project piplabs/story
1buf.build/gen/go/bufbuild/bufplugin/protocolbuffers/go v1.36.3-... h1:dS5ier+mttGuW+lRLP/eC1CKJm2Rg3rKWy9Iy0hroIU=
2buf.build/gen/go/bufbuild/bufplugin/protocolbuffers/go v1.36.3-.../go.mod h1:MYDFm9IHRP085R5Bis68mLc0mIqp5Q27Uk...
3buf.build/gen/go/bufbuild/protovalidate/protocolbuffers/go v1.36.3-...h1:cQZXKoQ+eB0kykzfJe80RP3nc+3PWbbBrUBm8XNYAQY=
4
5...
Listing 4: An excerpt from the lockfile (go.sum) of the project piplabs/story
3. Study of Lockfile Differences
Across ecosystems, lockfiles serve different purposes such as reproducible builds, package integrity verification, or performance improvement for project builds. Accordingly, different package managers make distinct design decisions when implementing the lockfile feature, depending on their priorities. It is our intuition that the different approaches to implementing lockfiles influence how developers use lockfiles. In this section, we systematically map the differences between the main lockfile designs, particularly the content of lockfiles and their lifecycles across various package managers. We then conduct a quantitative analysis of the number of projects that include lockfiles in version control within each ecosystem.
3.1. Research Questions
  1. RQ1:
    How do package managers generate, update, and structure lockfiles?
    Lockfiles have been independently designed and implemented across ecosystems, with some inspiration flowing from one to another. Lockfiles must include relevant, useful information without excessive verbosity, in order to keep both efficient project builds and high developer usability. This RQ highlights the diversity among package managers with respect to the generation, enforcement, and structure of lockfiles. Our study helps developers understand and adopt best practices for lockfile generation and management.
  2. RQ2:
    To what extent do open-source projects include lockfiles in version control?
    In this RQ, we quantitatively evaluate how developers of different ecosystems include lockfiles in version control. Including lockfiles in version control systems is important to achieve all benefits lockfiles have to offer, including deterministic builds (Goswami et al., 2020). We base our evaluation on lockfile commit data collected from open source GitHub projects created within the past 5 years (from 2019 to 2024). This quantitative analysis helps consolidate our observations in RQ1 about the usability and usefulness of lockfiles.
3.2. Methodology for RQ1
To answer the RQ1, we analyze the implementations of the lockfile feature in the source code of package managers, we review their documentation, and we experiment with sample projects that use each package manager.
Source code analysis: To understand the workflows of lockfile-related tasks for each package manager, we analyze their implementation within their source code repository. We begin by locating the main function that is invoked after executing an install or build command. We then analyze the functions called by this main function, manually tracing their execution flow. We further search the project using regular expressions for terms such as lockfile, locked, frozen, and dependency resolution. The sections of source code, analyzed for each package manager are summarized in Table 3. For instance, in the case of npm, we study the repository npm/cli, focusing on the source code for install and ci actions, dependency tree resolution, and lockfile validation.
Documentation: In addition to analyzing the source code of the 7 package managers, we review their official documentation. We focus on sections related to dependency resolution actions such as install or build, and study the available configurations or flags that can be used with these actions. We search for sections covering lockfiles, including lockfile-related recommendations and comparisons between lockfiles and dependency specification files. For example, as given in Table 3, for Poetry, we go through the section installing dependencies that explain lockfile usage, including installing without poetry.lock, installing with poetry.lock, and committing poetry.lock to version control. Similarly, for Pipenv, we analyze the commands and Pipfile & Pipfile.lock sections that cover general lockfile recommendations, examples, and security features.
Experiments with sample projects: To observe the dependency resolution behavior, with and without lockfiles, under different command-line options, we conduct experiments using sample projects for each package manager. First, we study the generation and validation of lockfiles under default configurations using the install or build command. Next, we experiment with various flags, such as --deploy or --locked, which trigger different lockfile validation behaviors. If we consider Gradle, as shown in Table 3, we build the project both with and without the --write-locks option and also with and without defining lock states in the dependency specification file. Further, we experiment with different locking modes strict, lenient, and default.
Table 3. Lockfile features for the 7 package managers under study. We study their source code, documentation, and experiment with sample projects.
Package manager Sections of source code Sections of documentation Commands run
npm CLI install and ci actions, dependency tree resolution (within arborist module), lockfile validation CLI commands npm install, npm ci
pnpm lockfile module, install action CLI commands - Manage dependencies pnpm i, pnpm i --frozen-lockfile, pnpm i --offline
Cargo install action, lockfile functions Build commands, manifest commands, Cargo.toml vs Cargo.lock cargo build, cargo build --locked, cargo build --offline, cargo build --frozen
Poetry installer class, locker class Installing dependencies poetry install, poetry sync
Pipenv command.py, install.py, locking.py, lock.py Pipenv commands, Pipfile & Pipfile.lock pipenv install, pipenv install --deploy, pipenv sync
Gradle locking package, dependencies package Locking versions gradle build, gradle build --write-locks (both with and without lock states in different modes; strict, lenient, and default)
Go modcmd, modload, modfetch, modget, mvs Managing dependencies Go Modules Reference go.mod file reference go mod init, go mod tidy, go get, go mod edit, go build
3.3. Methodology for RQ2
Next we aim to analyze lockfile usage across package managers, by selecting a sample of open source projects and collecting the lockfiles within them, if any.
Project selection: As the study subjects for RQ2, we collect popular open-source GitHub projects that use one of the 7 package managers we consider in the study. When collecting GitHub projects, we aim to exclude toy projects and only target mature, widely-used projects. To achieve this, we only consider repositories created between September 30, 2019, and September 30, 2024, that have at least 42 stars, 10 contributors, and 300 commits. To ensure that we only select actively maintained repositories, we further exclude the projects that have not been updated within the last 3 months. For npm CLI and pnpm, we select repositories where the primary language is JavaScript or TypeScript. For Pipenv and Poetry, we include Python projects. For Cargo, Gradle, and Go, we target Rust, Java, and Go projects, respectively.
Package manager identification: We filter the projects that use each package manager by selecting projects that have the corresponding dependency specification file in their root file directory. In JavaScript, there are multiple package managers that can share the same dependency specification file format and package registry. For example, the dependency specification file for both npm CLI and pnpm is package.json, as shown in Table 1. For Poetry as well, the dependency specification file pyproject.toml is compatible with multiple other package managers and packaging tools, such as pip and distutils. However, in Poetry dependency specification files, we can infer that a project uses Poetry by checking for section names such as [tool.poetry]. We exclude projects where pyproject.toml does not contain any Poetry specific sections. In Go, we consider only projects that use Go modules (introduced in Go 1.11 or later) and identify them by the presence of a go.mod file in the repository.
Lockfile identification: Similar to identifying the dependency specification file, we determine the presence of a lockfile by searching for the relevant lockfile files in the project directory. If a project has multiple lockfiles, we discard the project to prevent inconsistencies in the results and their interpretation. For example, in the JavaScript ecosystem, if a project contains both package-lock.json and pnpm-lock.json, we exclude it from the evaluation. We further disregard projects with lockfiles from package managers outside the selected 7 package managers, such as yarn.lock from Yarn. For Go, we check for the existence of the go.sum file. We also collect the project creation date and the earliest lockfile commit date. Using this information, we derive the number of projects that began including lockfiles in version control within the first six months of project creation, as well as the total count of projects that currently have lockfiles.
Excluding projects with no dependencies: For projects that use a package manager but do not declare any dependencies, lockfiles do not provide any added value. Therefore, we collect the number of direct dependencies specified in each dependency specification file and evaluate lockfile usage only for projects that have at least one declared direct dependency. In multi-module projects, if the main dependency specification file in the root folder does not specify any dependencies, we check the dependency specification files of sub-modules. If any sub-module declares third-party dependencies, we keep the project.
3.4. Results for RQ1: How do package managers generate, update, and structure lockfiles?
In this RQ, we first look at the fundamental differences in the lockfile content among the 7 selected package managers introduced in Table 1. Next we discuss the key differences in lockfile lifecycles including lockfile generation, updates and enforcement.
3.4.1. Lockfile Content
Package managers may include various information in lockfiles, incl. dependencies, checkums and additional details such as licenses. Figure 2 illustrates excerpts from lockfiles across the selected 7 package managers. Each of them contain an example representation of one dependency where the dependency versions, source links, checksums and indirect dependencies are highlighted in yellow, blue, green and pink respectively.
Table 5 presents the most common types of information recorded in lockfiles across different package managers.
Resolved package versions and checksums All of the studied package managers include resolved package versions in lockfiles. Checksums are also included in all lockfiles except in the Gradle lockfile. Gradle only records the groupId, artifactId, and version of the resolved dependencies which is a significant limitation.
Link to Source NPM and Cargo include a provenance link to source code. This is excellent for future rebuilding efforts (Lamb and Zacchiroli, 2022b). In terms of naming, npm CLI uses the field resolved for source links, while Cargo uses source. In Go, the source can be inferred from the module name declaration itself, as we see from the link to the GitHub repository for the module text in 11. Gradle, pnpm, Poetry, and Pipenv do not include a direct link to the source of the downloaded packages.
Indirect dependencies When recording indirect dependencies, lockfiles may keep the tree structure or flatten the list. Pipenv, Gradle, and Go, do not present indirect dependencies in a nested way. Go distinguishes indirect dependencies using the comment //indirect, whereas Pipenv and Gradle do not explicitly separate them from direct dependencies. npm, pnpm, Cargo, and Poetry provide a tree structure of depth 1 (as highlighted in pink in Figure 2).
Version ranges Cargo omits the version ranges of indirect dependencies (when listed under each parent dependency) unless multiple versions of the same dependency are required. When a version is included, Cargo specifies the exact resolved version. In contrast, npm, pnpm, and Poetry include version ranges for indirect dependencies under a parent dependency, reflecting the same constraints set in the dependency specification files. A better practice would be to display the exact resolved version.
Additional metadata Some package managers include additional details such as language requirements (e.g. Poetry, Pipenv), operating systems (e.g. npm, Cargo), or kernel compatibility (e.g. npm). Moreover, npm CLI includes in the lockfile extra metadata such as license details and funding information, directly copied from package.json (5). Although recording as much information as possible for each dependency may seem a good choice, it also makes the lockfiles lengthier and harder to be reviewed by human developers.
Summary of lockfile content
All the package managers in this study record both direct and indirect dependencies in lockfiles, with all except Gradle also including dependency checksums. This information is essential for verifying the integrity of a build or for reproducing a build. However, among package managers, there are notable differences in the additional metadata included and the way this information is structured, which impact the compactness and readability of lockfiles.
3.4.2. Lockfile Lifecycle
Table 4. Lockfile life cycles across package managers. All behaviors shown here are observed when invoking projects build under the default configuration. Special flags required to achieve the corresponding behavior are given within brackets.
Package manager Lockfile generation Lockfile based dependency resolution Lockfile enforcement - specification file mismatches
npm CLI by default by default optional (must use npm ci)
pnpm by default by default optional (only with --frozen-lockfile)
Cargo by default optional (only with --locked or --frozen flags) optional (only with --locked or --frozen flags)
Poetry by default by default by default
Pipenv by default optional (can use -deploy flag or sync action) optional (can use -deploy flag or sync action)
Gradle optional (only with –write-locks flag and specified lock states) by default by default
Go by default by default by default
1"node_modules/@eslint/eslintrc": { 2 "version": "2.1.4", 3 "resolved":"https://registry.npmjs", 4 "integrity":"sha512-269Z39MS6wVJts...", 5 "dev": true, 6 "dependencies": { "ajv": "^6.12.4", }, 7 "engines": { 8 "node": "^12.22.0" 9 }, 10 "funding": { 11 "url": "" 12 } 13} Listing 5: package-lock.json of the project asyncapi/modelina
1@vitejs/plugin-vue@5.0.4: 2 resolution:integrity:sha512-WS3hevEsz... 3 engines: {node: ^18.0.0 $|$ >=20.0.0} 4 peerDependencies: vite: ^5.0.0 Listing 6: pnpm-lock.yaml of the project csfive/composing-programs-zh
1[[package]] 2name = "Inflector" 3version = "0.11.4" 4 source = "registry+https://github.com..." 5checksum="fe438c63458706e0..." 6 dependencies = [ "lazy_static" ] Listing 7: Cargo.lock of the project leptos-rs/cargo-leptos
1[[package]] 2name = "jinja2" 3version = "3.1.4" 4description = "" 5optional = false 6python-versions =">=3.7" 7files = 8 [{file ="jinja.tar.gz", 9 hash="sha:2564a3aee7..."}] 10 [package.dependencies] MarkupSafe = ">=2.0" 11[package.extras] 12i18n = ["Babel (>=2.7)"] Listing 8: poetry.lock of the project mingrammer/diagrams
1"plexapi": { 2"hashes":["sha=2565af051..."], 3"index": "pypi", 4"markers": "python_version>=’3.9’", 5"version": "==4.16.1" 6}, Listing 9: Pipfile.lock of the project Drazzilb08/daps
1com.github.jknack:handlebars:4.3.1=testCompileClasspath,testRuntimeClasspath Listing 10: gradle.lockfile of the project datastax/fallout
1require ( 2 github.com/kr/text v0.2.0 //indirect 3) 4-------------------------------------- 5github.com/kr/text v0.2.0 h1:5Nx0Ya0ZqY2ygV366QzturHI13Jq95ApcVaJBhp 6github.com/kr/text v0.2.0/go.mod h1:eLer722TekiGuMkidMxCpM04lWEeraHUUmBw8l2 Listing 11: go.mod and go.sum (below the dashed line) of the project piplabs/story
Figure 2. Example dependency representations extracted from lockfiles (or a mod file) across 7 package managers. Dependency versions are highlighted in yellow, source links in blue, checksums in green, and indirect dependencies in pink. Some details are omitted for brevity and clarity.
All 7 package managers in this study include a function to generate a lockfile, either by default or when explicitly enabled during dependency resolution. During subsequent project builds, the presence of a lockfile may change dependency resolution. At each of these steps of lockfile generation, dependency resolution, and enforcement, different package managers implement different policies, which we detail below. A summary of the key differences are presented in Table 4.
Lockfile generation. Package managers take different approaches to generate lockfiles. All package managers except for Gradle generate a lockfile by default. For Gradle, a lockfile is generated only when the –write-locks flag is enabled and lock states are explicitly defined in the dependency specification file.
Lockfile based dependency resolution. When a lockfile is present at the beginning of a build, it influences the dependency resolution process. Lockfile based dependency resolution is the property of respecting the locked states when resolving dependencies as long as they do not conflict with the dependency specification file. In npm, pnpm, Poetry, Gradle and Go, the build process respects the lockfile by default and only includes the locked versions. Conversely, with the build or install command, Cargo and Pipenv allow resolving new dependencies regardless of the locked versions, ignoring the lockfile. Afterwards, Cargo and Pipenv silently update the lockfile during the build. To make Cargo respect an exiting lockfile, developers have to use --locked or --frozen flags and for Pipenv they can use the --deploy flag or the pipenv sync command.
Lockfile enforcement. Lockfile enforcement is the property of failing the build if the lockfile conflicts with the dependency specification file. Here, a conflict means a discrepancy in the checksum or the resolved version. In the presence of such a conflict, npm CLI and pnpm resolve versions compatible with the dependency specification file, overriding the lockfile silently (no enforcement). In contrast, some package managers fail the build in case of a conflict, thus enforcing the lockfile. This is the case for Poetry and Gradle where the lockfile is enforced by default, and an error is raised if conflicts are found. In npm, only npm ci command enforces the lockfile, and in pnpm, enforcement only happens with the --frozen-lockfile flag. In Cargo and Pipenv, the same flags required for lockfile based dependency resolution (--locked and --deploy) also achieve lockfile enforcement. For Go, lockfile enforcement happens by default because it does not have a separate developer maintained dependency specification file, and the checksums in go.sum are automatically validated.
Summary of lockfile lifecycles Except for Gradle, all package managers generate a lockfile by default but vary in how they consider locked versions. Cargo and Pipenv ignore them when resolving dependencies, whereas npm CLI and pnpm use them unless they conflict with the dependency specification file. Poetry enforces lockfiles strictly, stopping the build if conflicts are detected.
Findings from RQ1: Lockfiles address a fundamental problem of software engineering: knowing and verifying the exact versions of components required to build a software system. Cargo and Go lockfiles are feature complete. npm CLI lockfiles tend to be unnecessarily verbose due to excessive metadata. Gradle lockfiles are overly simplistic as they skip mandatory details such as checksums. Simple, streamlined lockfile processes such as Go’s are fool proof and recommended as blueprint for the future. Our findings are useful for developers to learn about lockfiles and for package manager maintainers to correctly design and implement effective lockfile features.
Table 5. The most common types of information included in lockfiles
Package manager Resolved package versions Package checksums Package source Direct/indirect differentiation Additional metadata
npm CLI $usym1F5F8 \usym$$\usym1F5F8$ Indirect dependencies (under each dependency), Dependency scopes, OS, engine, cpu, kernel requirements, Licenses, Funding information
[…]
2613
[…]
Cargo $usym1F5F8 \usym$$\usym1F5F8$ Indirect dependencies (under each dependency), Language, OS, engine requirements
[…]
2613
[…]
2613
[…]
2613
Language requirements
[…]
2613
[…]
2613
[…]
2613
Dependency scopes
[…]
3.5. Results for RQ2: To what extent do open-source projects include lockfiles in version control?
Table 6. GitHub projects analyzed to assess the presence of lockfiles in the version control system. The third column shows projects that started including lockfiles in version control within six months of project creation. The final column shows the number of projects that currently commit lockfiles.
Package manager No of projects with dependency specification files Projects with lockfiles within 06 months of creation Projects with lockfiles as of now
npm CLI 1916 43.4% (832) 53% (1021)
pnpm 18% (345) 36% (688)
Cargo 1089 54.7% (596) 70.9% (772)
Poetry 314 49.3% (155) 83.8% (263)
Pipenv 29 55.2% (16) 86.2% (25)
Gradle 323 0.3% (1) 0.9% (3)
Go 1188 92% (1095) 99.7% (1184)
Total 4859 62.6% (3040) 81.4% (3956)
Table 6 presents the total number of GitHub projects collected for each package manager, as well as the number of projects that publish their lockfile into the version control system. In the second column of the table, we report the number of projects that we collect, per the protocol introduced in subsection 3.3. The third column shows the number of projects that included lockfiles in their GitHub repository, in the early days of the project (within six months after project creation). The last column shows the projects that currently contain a lockfile.
In total, we collect 4859 projects that include a dependency specification file from one of the 7 package managers we consider. Within the first six months of project creation, 3040 (62.6%) of these projects began committing lockfiles. For example, 43.4% have committed a package-lock.json, and 18% have committed a pnpm-lock.json within six months of the project creation date. Over time, 916(18.8%) more projects across all package managers have included lockfiles in version control. As of March 28, 2025, 81.4% of these projects have committed lockfiles to version control.
The key finding from Table 6 is that the Go ecosystem has the highest level of lockfile support, with 99.7% of projects committing the lockfile to the repository. On the exact opposite, the state of lockfile usage in Gradle is poor with only 0.9% projects committing their lockfile. Below, we study these exceptional cases in detail, to check whether they reflect an anomaly in the data or whether they represent intentional design decisions.
For Go, most projects (92%) begin including go.sum within six months of project creation. The only exceptions are 4 projects: GoogleCloudPlatform/buildpacks, eolinker/apinto,
go-admin-team/go-admin, and shalb/cluster.dev. These projects have 13, 48, 29, and 35 direct dependencies in their main modules, respectively. The developers of the project, go-admin-team/
go-admin
committed the lockfile at the early stages of the project but later decided to remove it from version control. The reason for removal is not documented in the public repository. The projects GoogleCloudPlatform/buildpacks and eolinker/apinto have go.sum files for some modules but do not contain one for the main module.
GoogleCloudPlatform/buildpacks also uses the Go vendor folder which allows committing dependencies alongside the project, making them accessible at anytime. Vendoring loads dependencies from the local vendor folder instead of resolving them from remote repositories, thus supporting air-gapped builds. Yet, the go.sum file would still be useful to ensure the integrity of the dependencies properly.
In contrast, only 3 Gradle projects have a lockfile out of all 323 projects, and only one of them (apple/pkl) committed its lockfile within the first six months of project creation. The three Gradle projects that do commit lockfiles, apple/pkl, google/dwh-migration-tools, and datastax/fallout, are all multi-module projects, with 188, 461, and 53 dependencies across their modules, respectively. The google/dwh-migration-tools project’s issue tracker refers to an internal document that motivates the inclusion of the lockfile on Github. The project (apple/pkl) also makes use of finer-grained lockfile functionalities, such as selectively updating only certain dependency versions.
For the five other package managers, npm CLI, pnpm, Cargo, Poetry and Pipenv there is no clear pattern in lockfile usage. Poetry, Pipenv, and Cargo show relatively high adoption rates, with around 50% of projects committing lockfiles since the early stages and surpassing 70% as of now. In the JavaScript ecosystem, lockfiles are widely used but come from multiple package managers. Among them, npm CLI lockfiles appear to be the most common, more so than pnpm lockfiles. However, over time, the number of JavaScript projects committing pnpm lockfiles has doubled, suggesting a growing adoption of pnpm.
Quantitative data alone does not fully reveal the exact reasons why developers choose to commit or omit lockfiles. To complement this first quantitative analysis across ecosystems, research questions 3 and 4 will be qualitative developer interviews.
Findings from RQ2: Go has an almost 100% lockfile commit rate, in stark contrast to Gradle, where adoption is close to zero. Our empirical results show that generating a lockfile by default increases adoption. We recommend systematically including lockfiles in version control, as they help improve security and contribute to more efficient project builds.
4. Developer Interviews
Depending on the perceived usefulness and usability of lockfiles, and project specific requirements, developers may choose to actively use lockfiles or totally ignore them. In this section, we qualitatively analyze lockfile usage across projects in each ecosystem to identify differences in how developers perceive and use lockfiles. We also aim to discover developer expectations as well as the challenges they face when maintaining lockfiles.
4.1. Research Questions
  1. RQ3:
    What benefits do developers perceive in using lockfiles? Here, we study the benefits of lockfiles in terms of functionality, readability, and security. Moreover, we assess how developers’ perceptions of lockfiles vary across different package managers. Understanding the practical benefits of lockfiles helps clarify their role in secure and reliable software development.
  2. RQ4:
    What challenges do developers face when managing lockfiles? With this RQ, we study the effort required to learn about lockfile related functionalities, the maintenance burden, and situations where lockfiles fall short of meeting developer expectations. These findings are beneficial for package manager maintainers to improve their lockfile implementations for better usability and effectiveness.
4.2. Methodology for RQ3 and RQ4
To answer our 2 research questions on lockfile usage, we select a set of projects from each ecosystem, including some projects that publish their lockfile in the version control system and some projects that do not publish it. We then conduct semi-structured interviews with one of the top contributors of each project. In the following, we describe the methodology we follow to select the projects, contact developers, interview them and analyze their responses.
Participant selection: We aim to conduct interviews with developers who regularly contribute to open-source projects, have experience with managing dependencies, and maintain at least one project that uses one of the package managers under study.
To select projects that publish a lockfile, we proceed as follows. We first randomly select two projects per ecosystem from the list of projects that publish their lockfile, as identified in RQ2. We do not select Gradle projects that publish a lockfile, since we observe that this is an outlier behavior in the Java ecosystem (cf. subsection 3.5).
For package managers that have a considerable number of projects that do not publish a lockfile, we also select two projects without a lockfile. Therefore, for npm/pnpm, Cargo, and Poetry/Pipenv, we select a total of four projects each. We do not select Go projects that do not publish the go.sum, as this is an outlier behavior in the Go ecosystem.
For each selected project, we gather publicly available contact information for the top contributor and send an interview invitation. If contact information is unavailable, or we do not receive a response within a week, we continue iterating through our randomly selected projects, sending invitations until we reach our target of two or four participants per category, for a total of 14 responses. This approach allows us to select a diverse group of developers from various backgrounds and location, who have substantial expertise in dependency management and who contribute to projects in various application domains.
Participant demographics: In total, we sent 59 invitations and received 14 responses: two npm/pnpm developers, two Cargo developers, two Poetry/Pipenv and two Go developers who commit lockfiles. We also received responses from two npm/pnpm developers, two Cargo developers, and two Poetry/Pipenv developers who do not commit lockfiles. We also interview one developer (P1) we were acquainted with through personal connections. In total we interview 15 developers and Table 7 presents the demographics of these participants. They work in 12 different domains and have an average of 8 years of experience. For anonymity, we denote each participant with the letter ”P” followed by an integer.
Interviews: We conduct 20 minute online interviews with 15 selected developers. We record the interviews upon consent from the participants. One interview (P8) was conducted in person at the participant’s request and lasted approximately an hour; in this case, only written notes were taken. The participants were not provided any compensation for their participation.
We use two different versions of questions for the developers who commit lockfiles and for those who do not. Both sets of developers receive the same warm-up questions about their software development background, followed by a few interview preparation questions related to basic dependency management practices. The main interview questions are designed to address RQ3 and RQ4. Most questions are open-ended, encouraging a natural conversation and giving participants the opportunity to share experiences beyond direct responses. These questions cover topics such as the reasons for committing or not committing a lockfile, interactions with lockfiles (for developers who do not commit lockfiles, this pertained to analyzing locally created lockfiles), code review workflows, continuous integration and continuous delivery (CI/CD) practices, challenges encountered when using lockfiles, and suggestions for improving dependency management practices.
Interview response analysis: We use the Whisper speech recognition tool (Radford et al., 2023) to transcribe 15 interviews with available audio recordings. If audio recordings are not available, we refer to the written notes. Two authors carry out line-by-line coding for the transcribed interviews (or written notes). Consequently, the same two authors collaboratively create an affinity diagram, grouping codes into emerging themes inductively. The final affinity diagram consists of 20 sub-themes and 7 main themes related to general software development practices, lockfile benefits, developer expectations, and challenges.
Experience (Years) Package Manager Lockfile Usage Project Stats Project Domain
Stars Commits Contributors
P1 ¿ 20 npm pnpm Yes 3k 110,892 320 Website development
P2 2 - 5 Yes 24.4k 2552 221 Conversational AI
P3 2 - 5 Yes 9.6k 1240 136 Network monitoring
P4 5 - 10 No 2.7k 817 52 Bot framework
P5 10 - 20 No 219 476 21 Website deployment
P6 2 - 5 Cargo Yes 1.9k 1338 50 Untracked social media
P7 5 - 10 Yes 23.3k 2068 50 Network monitoring
P8 5 - 10 No 38.9k 8720 1221 Game engine
P9 2 - 5 No 639 1087 58 Bot API library
P10 5 - 10 Poetry Pipenv Yes 834 512 27 Home Assistant Integration
P11 5 - 10 Yes 391 413 16 Data modeling
P12 10 - 20 No 601 500 22 Network device connection
P13 2 - 5 No 708 529 13 Robotics
P14 10 - 20 Go Yes 670 324 12 CLI for ChatGPT
P15 10 - 20 Yes 948 1587 19 Email library
Table 7. Demographics of developers interviewed about their usage of lockfiles. The project domains are selected based on GitHub project descriptions. The experience indicates the number of years of experience of participants in programming in any language.
4.3. Results for RQ3: What benefits do developers perceive in using lockfiles?
One of the main topics we discuss in the interviews with developers who commit lockfiles is their motivation for doing so. From their responses we identify the benefits they expect from using lockfiles. Further questions about debugging, code review, and security practices also reveal the direct and indirect benefits lockfiles offer. We identify benefits across four main areas: build determinism, integrity, transparency, and debugging.
4.3.1. Build determinism benefit
One of the core functionalities of lockfiles is supporting build determinism over time by locking dependency versions. A majority of developers who commit lockfiles (6/9) mention build determinism as the main reason for committing lockfiles. For instance, both Cargo developers who commit lockfiles state that they view the lockfile as a snapshot, and they are confident that the build will be identical every time they use the same lockfile (”Like I’m 100% certain that it’s going to be identical no matter when I go back to it.”, P6). The Poetry developer, P10, mentions the guarantee of build determinism that lockfiles provide across different machines (”the major benefit I see is that my state is exactly the same when I pull down the code on a new computer. So it’s just great for getting everything into the perfect, the same exact state that I was working with before.”)
Resolving dependency versions deterministically is particularly important in CI/CD pipelines to ensure reliable test results across environments and consistent builds before releases. P1 comments that they enforce and trust the lockfile in their CI/CD workflow (”When we go through the pipeline in order to test or to run our tests, typically we use npm ci to install the dependencies by trusting the package lock.”) P10 mentions that they always make sure every collaborator uses lockfiles in their release pipelines (”So like in your release pipeline, you want to make sure you’re using your lockfile, right? You don’t want to just be updating your whole dependency tree. You want to be using your lockfile.”)
4.3.2. Integrity benefit
All package managers in this study, except Gradle, record dependency checksums in their lockfile. These checksums allow the package manager to verify, during subsequent installations, that the recorded checksums match with the dependencies being downloaded from the registry. This verification helps detect potential package tamperings. P10 emphasizes its importance, stating, ”If you don’t have a hash of your dependency, you don’t know what your dependency is. It’s just that simple, right? You’re flying blind.
All developers trust that the checksums recorded in lockfiles are automatically validated by the package manager, and therefore, they do not carry out any additional checks. P1 note that (”We trust the fact that the checksums are being validated, but we’d never really investigate that.” Similarly, P10 explained ”My understanding is Poetry already does validation for the hashes. So when I go and I download, when I go and I sync my repository on a new machine, if the hash of the package doesn’t match what’s in my lockfile, it will throw me a warning.”. Poetry keeps a hash of the dependency specification file content in the lockfile and performs an additional hash validation when checking whether the dependency specification file has been updated and the lockfile is outdated. Poetry developers, such as P10, appreciate this additional check, (”So it will say your lockfile’s out of date. And then it will make me regenerate my lockfile, which I appreciate.”).
Identifying package tampering through integrity verification requires that releases remain immutable. In Go, downloading packages directly from repositories like GitHub, where releases can be changed after publication, can lead to checksum failures if a package is updated under the same version. A Go developer described this as a trade-off between the convenience of releasing and the security. (”it’s good on the one hand that you will always, like from a security perspective, make sure that you will always get the code that is basically stored on the proxy. And if you don’t have that, something is wrong and the checksum will give you an error. But then again, it can cause issues if you like want to re-release the same version. And yeah, it’s like convenience versus security”, P15)
4.3.3. Transparency benefit
Lockfiles offer transparency into all dependencies resolved in a project, including indirect dependencies which are not visible in a developer maintained dependency specification file. When developers commit lockfiles, they can inspect changes to the dependency tree during code reviews and detect any unexpected or potentially malicious changes.
Both Cargo developers who commit lockfiles mention that they prefer to verify pull requests that modify the dependency tree by re-generating the lockfile locally to ensure the changes are reproducible. (”Since I prefer to double check, I clone their fork locally, and I run cargo update, so that I’m sure that the lockfile is reflected on the actual changes, and, you know, maybe they had something malicious, so it’s better to check.”, P7.) P6 confirms it stating ”Typically what I will do is, I’ll pull their change into a branch. I will recreate the change that they did to the cargo.toml and then see if my lockfile changed in the exact same way.
Lockfiles, by exposing indirect dependencies, help developers monitor and manage dependency growth over time. P15 recall a time where they advised a collaborator to remove an unnecessary dependency ”So I said, okay, let’s not just import this whole dependency just for one Go file. Let’s just import the code over because it’s probably not changed anyways. And yeah, so usually I check the dependencies and basically question is this dependency needed or not”.
4.3.4. Debugging benefit
Almost all developers have experienced breaking dependency updates and semantic versioning (semver) violations in their projects. Developers who commit lockfiles use them to help in debugging breaking changes. P2 recollect how they copied a previous version of a lockfile to resolve a breaking update, ”That was one time I literally manually copy-pasted. I went exactly to where that package was resolved before, and it was working, and I literally replaced it”.
Another advantage of using locked versions is that developers can explicitly control when to allow dependency updates, preventing unexpected breakages. P11 mention that they update the lockfile only when they are prepared to handle potential breaking changes. (“From time to time, I run poetry updates with the intention of getting the latest and being in a mindset where I now have the surplus time and energy to handle breaking changes or issues that could arise from using the latest versions.”) P7 mention how they have to enforce the lockfile to prevent a dependency from getting updated because it does not follow the semantic versioning guidelines. (”So basically, it was the same problem. Semantic versioning was not respected by them. So I updated the README, and saying that if they download from cargo, they should use cargo install –locked. Because in that way, I’m sure that they’re using the dependency in the lockfile.”).
In Cargo, there is the yanked release issue, when dependency is no longer available by version, but remain available by checksum. The Cargo dependency resolver ignores yanked versions unless they are specified in the lockfile, occasionally causing the build to fail. P6 discuss how they overcome the yanked version problem in Cargo through lockfiles, ”If you have the lockfile, you can always build it, even if it’s yanked. But if you don’t have a lockfile and it gets yanked, you cannot under any circumstances build it.”.
4.3.5. Security benefit
Another key point raised by the developers is that lockfiles are indirectly used by other security scanning tools. For instance, P1 mention using the Black Duck software component analysis tool, which relies on lockfiles (”Black Duck uses a list of vulnerabilities based on the package names, on their versions, and reports on a regular basis on our products to know if one of the dependencies of this product is having a known vulnerability. So that’s one thing which leverages the package-lock.json”). When a lockfile is present, dependency management bots such as Dependabot can also detect and flag vulnerabilities within indirect dependencies (”It’s definitely beneficial to have a lockfile for the purposes of Dependabot. Yeah. Dependabot is what I rely on heavily for making sure I’m dealing with supply chain security properly… So Dependabot will scan the lockfile in order to determine the whole dependency tree, and then it wll warn you for CVEs even within transitive dependencies”, P10).
Findings from RQ3: Lockfiles offer benefits across various dimensions of software development, including build determinism, integrity, transparency, debugging, and security. Developers who commit lockfiles, regardless of the package manager, are generally aware of these benefits. In each ecosystem, many developers have experienced unique software development problems where lockfiles is the solution. Overall, developers who use lockfiles see them as an essential component for effective dependency management.
4.4. Results for RQ4: What challenges do developers face when managing lockfiles?
As part of our interview protocol, we ask developers who do not commit lockfiles about their reasons, and those who do commit lockfiles about the challenges they face when using them. From their responses, we identify several common challenges, including transitive locking with libraries, managing dependency updates, limited readability, cache invalidation, and the learning curve associated with lockfile practices.
4.4.1. Indirect dependency locking challenge
The main reason mentioned by four out of six developers who do not commit lockfiles is that their project is a library rather than an application or executable. These projects are in Python, Javascript or Rust and are in various application domains. The challenge they face is that even if they lock the versions of dependencies in their project, they cannot publish a lockfile to the package registry, and downstream dependents could not enforce the lockfile for their library. As a result, they choose not to enforce a lockfile in their development environment, as doing so would give them a false sense of security. (”So basically it is better not to commit a lockfile because then you kind of run into these issues faster. And, uh, yeah, you don’t have this false sense of security where you think it’s all locked, but in reality, everybody who uses it doesn’t lock it.”, P4). Following the same reasoning, P8 explain that they prioritize the experience of the users of their library over their own developer experience, and as a result, they do not commit lockfiles.
Similarly, some developers view determinism as a concern of downstream users who depend on their library. (”You could make your end thing deterministic, but as a library author, for me, it’s, that’s not really my problem.” P12).
P9 do not currently commit lockfiles and they mention that they would reconsider their decision based on the updated Cargo guidelines introduced in 2023. They refer to the StackOverflow thread Should Cargo.lock be committed when the crate is both a Rust library and an executable. (”This was the previous recommendation on StackOverflow, but has since changed. I have not revisited this choice, but if I were to start a new project today, I would follow the latest recommendations and defaults”, P9.)
4.4.2. Slowing down dependency updates
A second challenge developers face with lockfiles is that version locking can slow down frequent dependency updates. Almost all developers use bots such as Renovate and Dependabot to manage dependency updates, which also handle automatic lockfile updates. However, they are concerned about the lag associated with updating dependencies through bot-generated pull requests. P8, in particular, mention that they prioritize frequent dependency updates over deterministic builds to provide a better experience for their library users.
P11 mention that, with Poetry, if a lockfile references a version of a dependency that has been yanked, it results in an error, unlike Cargo. However, if they do not have a lockfile, Poetry will resolve a newer version, without raising an error. They explain the issue with an example. ”Now, if you, at 9 a.m. in the morning, let’s say a library pushes a new version. At 9.30, you create a lockfile with that version locked. And at 10 a.m., they delete that version from PyPI. Well, then anybody who syncs using that lockfile hits a big issue. Because that version is no longer available”. At the same time, they acknowledge that this behavior can sometimes be helpful. ”it reports it as yanked, and therefore doesn’t install it. So you get an early error, which in some situations is desirable”.
In npm, the current state of project dependencies is referred to as the ideal state. For projects without a lockfile, npm CLI always resolves dependencies to this ideal state at the time of installation. P5 prefer resolving this ideal state rather than relying on a potentially outdated lockfile. ”Because you’re always resolving the ideal tree, like at a given moment, that tends to be a better resolution than whatever state your lock file is accumulated over six months or whatever”. They also believe that continuously resolving the ideal tree helps fixing bugs and vulnerabilities more effectively. ”Oftentimes, either you’ll just miss the defect because it will be introduced for a short period of time and then got published with a fix. And so you just won’t even notice it just because you’re always resolving the ideal tree”.
4.4.3. Human readability challenge
All five npm/pnpm developers mention that they find the lockfiles too long, verbose and hard to read. For example, P1 state ”it’s quite wild what you can find inside the package lock. So it’s not really readable”, whereas P2 mention how this verbosity can complicate the code reviews and continuous integration, ”Git is kind of overloading and taking up a bunch of memory. And is even having trouble showing them all”. P5 also comment on the human readability of lockfiles, ”save us from the insanity that these machine generated lock files, which are just, you know, they’re human readable, but they’re not because they’re so big”. Poetry/Pipenv developers raise the same issue where P10 explain, ”it’s like a huge JSON file. So it’s kind of annoying to review the diff. And I don’t get a lot of value out of reviewing the diff, necessarily.”. P11 add that they also find it difficult to code review the lockfile, ”I will be much more likely to review any changes to a pyproject.toml or, you know, direct dependency file. So I’d say it’s to some degree a weakness of lockfiles that they are so big”.
On the other hand, none of the Cargo or Go developers raise issues related to the readability. Go developer P15 specifically mention they find reviewing go.mod files very easy. ”But usually my Go mod files are very like, they do not go over one page of my screen”. P12 who develop in Python and do not use lockfiles also comment that they get a better developer experience with Go, ”It’s basically almost the same because go.mod is also kind of very minimalistic, So I guess you get all of the benefits of like, you know, pip lockfile or poetry lockfile or whatever with go without having to think about it.” This confirms the fact that, depending on the lockfile implementation details, such as the content included in lockfiles, developers either effectively use them for manual code reviews or tend to overlook them entirely.
When asked about the information they consider during code reviews of lockfiles, the majority of developers who commit lockfiles mention that they focus on the name, URL, version of the package, and the number of indirect dependencies it includes. A consistent observation we make is that none of the developers pay attention to the checksum of the dependencies in the lockfile. (I’ve never actually checked checksums, P7). Following the same reasoning, Go developers recommend the Go’s approach of having checksums in a separate file. (”To be honest, I really don’t care where the checksum is stored. It makes it clean to have it separated because I only have to put it into the Go mod file and then the checksum will be generated automatically in the some file. So that’s like having a look at your Go mod file is much cleaner., P15 ”).
4.4.4. Cache challenge
A common issue faced by all developers is cache invalidation. P8 remark that ”Cargo caching mechanisms are buggy,” whereas P1 mention that they have to recreate the lockfile to address cache invalidation problems, ”I would say most of the time what people do is they delete the node modules folder, they delete the package lock, and then they retry and install”. Following up on cache invalidation, P2 mention that they occasionally run into operating system specific issues which cannot be solved by simply deleting the cache (”Just working across different operating systems, um, you know, also on macOS. I’ve had issues where deleting, like doing that same process, deleting the node modules folder, created issues with npm and then nothing would install.”). Go developers also face cache related issues with Go proxy when packages get re-released under the same version (”Go proxy already had that basically broken release downloaded and stored in its proxy cache. So people were like opening issues that they had checksum issues”, P15).
4.4.5. Learning challenge
Another point raised by developers is the learning curve of lockfile related practices across different package managers. P14 compare Gradle and Go noting that they find it easier to learn and understand dependency resolution practices in Go, ”it’s very, very simple. It doesn’t have much of a learning curve, such as like Maven or, Gradle”. P12 also comment on this ”I think Go is just an easier, it’s just easier. It works better. And, and, you know, it’s a newer language and lessons learned”.
npm developers often mention that the logic behind dependency resolution and lockfile creation processes are unclear (”maybe an understanding of how it constructs the lockfile. Cause it’s still kind of a mystery to me”, P2). Many also comment on the unhelpfulness of error messages when debugging lockfile related issues (”The error message that you get is really cryptic.”, P1). They note that resolving these issues requires an in-depth understanding of the underlying processes, and the debugging itself is extremely time consuming (”Some of these issues I’ve had with it are just so hard to debug until I do some very thorough digging of what’s going on”, P2). P15 compare npm CLI with Go and explain that the minimalism makes the dependency resolution process in Go more transparent and easier to understand. ”You just say npm install and like a wall of text goes down and you have a gazillion packages installed and you don’t really know what happens and why those are installed. Compare that to Go, which is, in my opinion, very open and very clear on what dependencies you are using.
Findings from RQ4: Developers who do not publish their lockfile in version control, do so for clear reasons. These reasons include avoiding lockfiles when developing libraries, readability challenges due to the verbosity of lockfiles, or avoiding delays in dependency updates. Overall, developers agree on the fact that lockfile challenges are better managed in Go than in JavaScript or Python.
5. Informing the Future of Package Managers
In our study, we have analyzed package managers and lockfiles from diverse perspectives, reading code and documentation, comparing features and contents, and interviewing users. This gives us a unique perspective for reflection on the principled engineering of lockfiles for a healthy software supply chain.
[Recommendation #1] We recommend committing lockfiles for all projects.
Commiting lockfiles allow for distributed integrity checks and deterministic builds. This holds for libraries and applications, as library developers face the same security risks and challenges of collaborative development as application developers. To avoid unexpected issues with in-range breakages that downstream users may face, application developers should be able to perform dependency resolution in different ways, with or without enforcing locked versions.
[Recommendation #2] We recommend future package managers to prioritize developer experience.
Having multiple configuration options for lockfiles gives developers more flexibility and control. Yet, it can also discourage them from using the basic functionalities of lockfiles due to the added complexity. For less experienced developers, setting up appropriate configuration can be particularly confusing. For instance, in Gradle, developers have to enable a flag, update the dependency specification file with lock states, and select an appropriate mode to generate a lockfile. On the other hand, sensible defaults and strict enforcement simplifies the process of using lockfiles and makes it easier to follow best practices. In Go, the complexity of lockfiles is remarkably reduced to a bare minimal developer intervention.
[Recommendation #3] We recommend lockfiles to include only essential content.
When designing lockfiles, package managers should avoid duplicating information already present in the dependency specification file. They should also exclude data that is irrelevant to dependency locking. In npm, for example, including data such as licenses and funding information in the lockfile adds little to no value while unnecessarily increasing code reviewing cost. At the other extreme, Gradle presents a serious lockfile anti-pattern by omitting essential details such as checksums. We recommend lockfiles to include the following mandatory details: resolved package versions, checksums, URLs for downloaded packages, and a clear distinction between direct and indirect dependencies. Any additional metadata should be kept as brief as possible, and ideally included only for reproducibility purposes, such as OS, engine, and language versions. We also encourage separating checksums into a separate file, similar to the approach of Go, to improve the human readability of lockfiles.
[Recommendation #4] We command package managers to generate lockfiles by default.
Lockfile generation should be enabled by default as lockfiles provide unique software engineering benefits, including integrity verification, reproducibility, and performance improvements, all no cost if enabled by default. We claim that dependency resolution should respect the locked versions by default and should never silently update the lockfile to prevent any unexpected build changes. This is essential to detect resolution based software supply chain attacks. As example to follow, Poetry respects locked versions and warns when the lockfile is out of date without silently updating it, we observed that developers who care about project security do appreciate this behavior. On the other end, when the lockfile is not generated by default, as with Gradle, developers seldom use them.
6. Related work
To the best of our knowledge, lockfiles have not been studied in detail in the literature, a surprising fact given their importance in the field. Some studies occasionally mention them in the broader context of dependency management.
6.1. Dependency Resolution and Package Managers
One of the key responsibilities of a package manager is resolving compatible versions for all the dependencies in a project. Zacchiroli et al. (Abate et al., 2020) review the usage of dependency resolving algorithms, including specialized constraint solving algorithms as well as ad-hoc dependency tree based algorithms. They conduct a census of dependency resolution mechanisms across 18 package managers, using version locking as one of the criteria for comparison. Pinckney et al. introduce a unified framework built on top of Max-SMT solvers to resolve dependencies more systematically, moving beyond ad-hoc algorithms (Pinckney et al., 2023).
Several studies focus on dependency conflict issues in different ecosystems. For example, (Wang et al., 2020) identifies factors that lead to conflicts, whereas (Wang et al., 2023) highlights issues related to inefficiency and excessive resource usage by the dependency resolution strategies in Python. (Wang et al., 2019) introduces a tool to generate crash stack traces for dependency conflicts in Java. Patra et al. (Patra et al., 2018) analyze dependency conflicts in JavaScript that arise from namespace collisions. (Wang et al., 2025) presents an in-depth study of conflicts between peer dependencies in npm. Similarly, in the Go ecosystem, Wang et al. (Wang et al., 2021) study conflicts caused by the coexistence of two library referencing modes: Go PATHs and Go modules.
Liu et al. (Liu et al., 2022) analyze vulnerability propagation within dependency trees by applying npm-specific dependency resolution rules, rather than relying on simple reachability analysis for indirect dependencies. They recommend managing dependencies with lockfiles. Along similar lines, He et al. (He et al., 2025) study the effect of dependency pinning in npm projects. They find that local pinning leads to more security vulnerabilities due to bloated and outdated dependencies. They also suggest that the risk of malicious package updates can be reduced when a small set of core dependencies pin their versions and keep them updated regularly.
These previous studies analyze dependency resolution mechanisms either directly or in the context of vulnerability detection. Our work complements these works with an analysis focused on the impact of locking of dependency resolution across package managers.
6.2. Dependency Management Practices in Open Source Software
Several empirical studies provide insights into the dependency management practices of open source software projects. Bogart et al. (Bogart et al., 2021) discuss practices related to breaking changes across different package managers, and highlight the use of lockfiles to lock versions of indirect dependencies as a means to stabilize projects. Breaking updates have been well studied in the literature (Ochoa et al., 2022; Reyes et al., 2024; Frank et al., 2024). Venturini et al. (Venturini et al., 2023) quantitatively evaluate the impact of breaking updates on dependent packages in the npm ecosystem. They also note the advantages of using lockfiles to track all resolved packages and avoid in-range breaking changes.
Schueller et al. (Schueller et al., 2022) curate data from the Rust ecosystem over eight years, capturing developer activity, library dependencies, and usage trends. Another empirical analysis in Rust (Li et al., 2023) reveals that 46% of packages adopted yanked releases and the proportion of yanked releases keeps increasing over the years. In Cargo, such yanked releases can be resolved only if a lockfile is present.
There is a body of work that analyzes security practices and threats associated with dependency management. Gu et al. (Gu et al., 2023) conduct a systematic analysis of security threats in package registries; Maven, PyPI, npm, Cargo, NuGet, and Go. Rahman et al. (Rahman et al., 2025) propose a metric to evaluate vulnerabilities in dependencies based on the time it takes for a fix to be applied. Bos (Bos, 2023) reviews supply chain attacks and discusses a type of attack that originates from lockfile tamperings. One such tampering is adding a duplicate entry to the lockfile with a modified package and its new checksum. As most lockfiles are tediously long for thorough review, such tampering goes undetected leading to supply chain attacks. Vaidya et al. (Vaidya et al., 2019) characterize attacks in npm and PyPI ecosystems. Their findings suggest that with proper tooling and practices it is possible to mitigate these attacks to a great extent. The authors discuss the npm CLI lockfile npm-shrinkwrap that allows indirect dependency locking. They mention that npm-shrinkwrap can prevent future attacks from new packages through dependency locking. Kabir et al. (Kabir et al., 2022) study security best practices in software development and consider committing lockfiles as one of the best practices.
Most of the related work on dependency management practices discusses lockfiles only briefly. Our contribution is unique by being entirely focused on lockfiles, and providing novel software engineering knowledge on their usage and adoption.
6.3. Adoption of Dependency Management Tools within Development Teams
Previous works have also used combined qualitative and quantitative methods to learn about developers’ adoption of dependency management tools. Schorlemmer et al. (Schorlemmer et al., 2024) study the adoption of software signing in four package registries; Maven, PyPI, DockerHub and Huggingface. Kalu et al. (Kalu et al., 2024) conduct a follow-up interview study to understand software signing in practice. They identify that strict signature rules increase the quantity of signatures, and that registry policies have an impact on the developer decision to adopt proper signing.
Saroar et al. (Saroar and Nayebi, 2023) survey 90 developers to understand how they use, choose, and troubleshoot GitHub Actions. Holtgrave et al. (Holtgrave et al., 2025) analyze different attack types that can be performed by impersonating GitHub user accounts. Bifolco et al. (Bifolco et al., 2024) assess the accuracy of the GitHub dependency graph in Java and Python projects. Bifolco et al. use lockfiles as one source of truth to assess the GitHub dependency graph.
Dependency management bots such as dependabot, renovate and Greenkeeper automate the process of updating dependency specification files and lockfiles when new dependency versions are available. Alfadel et al. (Alfadel et al., 2021) and Mohayeji et al. (Mohayeji et al., 2023) evaluate how developers respond to security updates suggested by dependabot. Rombaut et al. (Rombaut et al., 2023) study whether Greenkeeper reduces developer effort or introduces unnecessary workload. They mention lockfiles as a way to overcome in-range breaking changes, and also briefly mention about the challenge library developers face for sharing lockfiles. Our qualitative analysis confirms that developers do use bots to maintain dependency specification files and lockfiles up to date.
Our novel findings on lockfiles, their contents and their adoption in practice contributes to the broader body of knowledge about the state of practice regarding the state of the art of dependency management tools.
7. Conclusion
We have conducted a comprehensive analysis of lockfiles across 7 package managers, analyzing their documentation and implementation, and collecting quantitative data from open-source GitHub projects and qualitative data from developer interviews. To our knowledge, this is the first in-depth study conducted specifically on lockfiles. We systematically determine how different package managers follow different approaches to generate and enforce lockfiles, and these decisions ultimately result in varied levels of developer satisfaction. We further identify the reasons why developers find lockfiles useful: to achieve deterministic builds, to perform integrity checks, to simplify debugging. We also unveil the unique challenges developers face when using lockfiles: readability issues, indirect dependency locking limitations, and delayed dependency updates. Based on these observations, we provide four recommendations: always generate a lockfile, always commit the lockfile, limit lockfile content to essential metadata, and consider developer experience.
Today, only the Go ecosystem systematically generates and adds lockfiles in version control system. The Go lockfile is particularly readable, concise and clearly separates the list of dependencies from their checksums, making them easier to understand for developers. On the other hand, lockfiles are almost absent from the Java ecosystem. Out of the 7 package managers we study, only Gradle does not generate a lockfile by default. Meanwhile, Maven, the other major package manager for Java does not have a lockfile at all. We command the Maven community add this feature and learn from the best practices to design an informative and usable lockfile.
As future work, we shall experiment with ecosystem locking to capture the overall state of the package registry, instead of locking individual packages. We shall design and implement support for developers to toggle indirect lockfile enforcement and to allow library authors to publish a lockfile to the package registry, for giving downstream users the flexibility to choose between locking dependencies or allowing floating versions. Clearly, the destiny of software integrity lies on lockfiles.
Acknowledgments
We thank all the interview participants for dedicating their time for this work. We acknowledge the valuable insights provided by Ruy Adorno, a software engineer at vlt. We also thank Ian Arawjo for feedback on early versions of this paper. This work is supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the IVADO consortium.
Appendix A Lockfile Related Functions
npm CLI: npm CLI generates a lockfile by default with the npm install action when executed without additional flags. When no lockfile is available on the first run, npm CLI uses the current state of installed dependencies on disk within the node_modules folder as a starting point. Then, according to the dependency specification files, i.e., package.json, it resolves direct and indirect dependencies by connecting to the npm registry. On the other hand, if a lockfile exists already, npm CLI combines the lockfile and the dependency specification file content to construct the dependency tree. Therefore, the dependency versions resolved would be the same as the locked versions in the lockfile if there are no conflicts between the dependency specification file and the lockfile. If a conflict is detected, npm CLI resolves the dependencies according to the dependency specification file and automatically updates the lockfile. For example, in 1, if the version of the dependency apidevtools/json-schema-ref-parser is updated to ^11.2.0 from ^11.1.0, it conflicts with the version 11.1.1 locked in the lockfile. In such cases, npm CLI resolves a newer available version that satisfies the given version range and updates the lockfile accordingly. When running the npm ci action, the lockfile is validated strictly. This means that npm CLI compares the desired state of dependencies in the dependency specification file, with the dependency tree derived from the lockfile, and raises an error if any mismatch is detected.
pnpm: The pnpm install action functions similar to the npm install action. It creates a lockfile by default when a lockfile does not exist. In case the lockfile exists, it considers the versions locked in the lockfile as long as there are no mismatches with the dependency specification file, or updates the lockfile in case of mismatches. The equivalent behavior to the npm ci action can be achieved with the --frozen-lockfile flag in pnpm.
Cargo: Cargo generates a lockfile by default with cargo build action. However, unlike npm CLI and pnpm, the Cargo build action does not consider the lockfile when resolving dependencies even when one exists. If no lockfile exists or changes to dependencies are required, Cargo automatically writes to the lockfile. To enforce strict use of the locked versions and to disallow writing to the lockfile, developers must enable the --locked or --frozen flags. Compared to the --frozen flag in pnpm, Cargo --frozen flag not only enforces locked versions but also operates in offline mode, supporting air-gapped builds.
Poetry: If no lockfile exists when executing the install or sync actions, Poetry resolves the dependencies and generates a new lockfile by default. If a lockfile exists, Poetry relies on it during the install action, and does not modify it under default configurations. Instead, it raises an error if the lockfile needs to be updated to align with the dependency specification file. During the install action, Poetry compares the content hash of the dependency specification file with the hash recorded in the lockfile’s metadata section. Matching hashes indicates that the dependency specification file has not been updated. However, if the hashes differ due to changes in the dependency specification file, the install process terminates, recommending that the lockfile be updated first by running the poetry update action. The update action resolves the latest versions of dependencies and updates the lockfile accordingly. Similar to npm ci and cargo build –locked actions, Poetry has the poetry sync action that ensures only the dependencies and versions specified in the lockfile are resolved and installed in the project environment.
Pipenv: Pipenv follows a process somewhat similar to Poetry for its install, update, and sync actions. However, under default conditions, Pipenv does not throw an error if the dependency specification file content hash differs from the recorded hash in the lockfile when executing the install action. By default, it also does not strictly rely on the lockfile when resolving dependencies. If new versions are available under the version range defined in the dependency specification file, it will resolve the newer versions irrespective of the locked versions in the lockfile. With Pipenv, the lockfile is only enforced (only the locked dependency versions are considered), and a hash mismatch triggers an error, when the --deploy flag is enabled during the install action. The sync action in Pipenv behaves similarly to Poetry’s, and also enforces locked versions. However, unlike Poetry’s sync action, Pipenv throws an error if a lockfile is missing when the sync action is executed.
Gradle: Gradle takes a distinct approach to generating and enforcing lockfiles compared to other package managers. By default, the Gradle build action neither considers an existing lockfile nor creates a new one when a lockfile does not exist. A lockfile is generated only when the --write-locks flag is enabled. However, even with this flag enabled, a lockfile is not created unless lock states are explicitly defined in the dependency specification file. Developers must specify which groups of dependencies they wish to lock in the dependency specification file. Once these lock states are defined and the --write-locks flag is enabled, Gradle will either create a new lockfile or update an existing one. If a lockfile already exists, Gradle provides options for enforcing the lockfile. In the default and strict modes, Gradle validates the lockfile and throws errors if there are any dependency mismatches. In the lenient mode, however, mismatches with the lockfile do not trigger an error.
Go: Go uses a deterministic dependency resolution algorithm called Minimal Version Selection (MVS), and two files: one called go.mod that works as both a dependency specification file and a lockfile, designed to be machine-written and human-readable, and another called go.sum that stores package checksums. When creating a Go module, the go mod init command generates a go.mod file that contains the Go version and the module path. As they implement their modules, developers import dependencies within the source files that require them. Subsequently, running go mod tidy or any module-building command, such as go build, adds module dependencies resolved according to the import statements to go.mod and generates the go.sum. For additional edits developers can use the go get or go mod edit commands.